OpenStudy (anonymous):

simplify (12)/(2sqrt(12x))

jigglypuff314 (jigglypuff314):

\[\frac{ 12 }{ 2\sqrt{12x} }\rightarrow \frac{ 6 }{ \sqrt{12x} }\] reduce as much as you can first, then \[\frac{ 6 }{ \sqrt{12x} }= \frac{ 6 }{ \sqrt{4*3*x} } = \frac{ 6 }{ \sqrt{4}*\sqrt{3x} } = \frac{ 6 }{ 2\sqrt{3x} } = \frac{ 3 }{ \sqrt{3x} }\] then if you don't want a square root in the denominator \[\frac{ 3 }{ \sqrt{3x} }*\frac{ \sqrt{3x} }{ \sqrt{3x} } = \frac{ 3\sqrt{3x} }{ 3x } = \frac{ \sqrt{3x} }{ x }\]